OpenStudy (anonymous):

Trying to figure out 5(2n-3) squared, calculator says 25 trying to figure out how Thanks

OpenStudy (anonymous):

= 25(2n-3)^2 = 25(4n^2 -12n +9) = 100n^2 - 300n + 225